# Contrastive Retrieval Heads Improve Attention-Based Re-Ranking

### v1

@@ 5.2 Results
Re-ranking performance.
Table 1 presents the re-ranking results on the BEIR benchmark using the granite-embedding-30m-english retriever baseline. On average, CoRe-R achieves the best performance across all four models, consistently outperforming both ICR and QR-R. Aggregating attention from all heads in ICR yields the weakest results, in some cases falling below the retriever baseline for Mistral 7B and Phi-4. In contrast, CoRe-R delivers substantial improvements over the retriever baseline and ICR on every model. Compared to QR-R, CoRe-R shows the largest gains on Mistral 7B with average improvements of +2.0+2.0 points over QR-R and +3.0+3.0 points over ICR, and on Llama-3.1 8B with +0.8+0.8 and +1.9+1.9 points respectively. While QR-R remains competitive on the stronger Phi-4 model, its performance fluctuates across datasets, whereas CoRe-R provides more consistent and robust improvements. These findings demonstrate that CoRe-R effectively isolates meaningful attention heads across models and datasets, establishing it as a reliable state-of-the-art list-wise re-ranker.
Table 1: nDCG@10 on the BEIR benchmark.
@@
TREC-COVID 63.1 72.1 73.1 73.8 76.8 75.7 75.7 75.2 76.8 76.0 70.4 70.3 71.8
NFCorpus 33.7 33.1 35.0 35.2 35.0 36.4 36.7 31.6 35.5 35.6 31.8 32.5 33.9
DBPedia 36.0 36.0 36.8 37.5 38.7 39.0 39.4 36.9 39.4 39.6 35.6 35.9 36.8
SciFact 71.3 71.3 71.9 74.7 74.7 75.9 75.1 71.7 74.2 74.7 74.6 74.7 75.2
SciDocs 22.5 16.6 17.3 19.5 18.5 19.7 20.2 17.8 21.0 21.1 18.4 20.3 20.0
FiQA 36.9 35.8 40.3 41.4 41.3 44.0 44.1 37.2 43.8 44.1 39.7 41.6 42.4
NQ 51.6 55.4 56.5 58.1 60.8 63.0 63.2 57.4 63.3 63.6 57.4 57.9 60.7
FEVER 85.5 87.5 87.1 88.7 89.2 87.4 88.4 89.1 89.6 89.6 88.0 87.5 86.2
Climate-FEVER 30.3 21.6 21.1 23.3 22.5 22.9 22.7 22.2 24.5 24.3 20.5 21.2 21.0
HotpotQA 62.9 71.7 71.6 73.4 73.7 73.6 73.8 73.3 74.2 74.5 73.6 72.8 72.6
Touche 24.0 23.3 26.4 26.7 26.1 26.3 26.4 21.8 25.4 25.7 20.5 24.6 24.8
MSMARCO 30.7 30.1 32.1 31.7 32.9 33.9 34.9 30.6 34.6 34.8 29.6 31.2 32.5
Quora 86.7 81.1 71.5 85.2 78.7 74.6 82.3 74.2 76.5 80.2 83.1 79.8 75.1
ArguAna 56.4 45.0 51.1 52.7 42.5 54.2 55.0 41.7 52.2 52.7 52.3 56.1 57.3
CQADupstack 44.3 38.7 40.8 41.6 41.7 43.2 43.9 39.4 44.0 44.3 41.4 42.3 43.4
Average 49.1 47.9 48.9 50.9 50.2 51.3 52.1 48.0 51.7 52.1 49.1 50.0 50.3
Table 2: nDCG@10 on the MLDR datasets.
@@
Cross-lingual Generalization.
Table 2 reports the nDCG@10 on the MLDR datasets across three language models. We exclude Phi-4 as the model is not intended to support multilingual use (Microsoft Research, 2024). To ensure compatibility, we evaluate only languages that are supported by all three models. With the same set of CoRe heads detected using a subset of the NQ dataset, CoRe-R shows major improvement in the re-ranking accuracy over all baselines. For Mistral 7B, both ICR and QR-R underperform the retriever baseline on average, while CoRe-R achieves an average improvement of 2.8 points. For Llama-3.1 8B and Granite-3.2 8B, CoRe-R attains the best nDCG@10 scores on all languages with an exception of German for Granite. Overall, CoRe-R demonstrates consistent cross-lingual gains beyond English, underscoring the strong generalization ability of CoRe heads.
@@
Figure 3: Average nDCG@10 on BEIR benchmark using the attention signal from different number of top retrieval heads.
Less Is More: Re-Ranking with Few CoRe Heads.
Figure 3 illustrates the re-ranking performance as a function of the number of retrieval heads across three models. Across all settings, CoRe-R consistently outperforms ICR, demonstrating the effectiveness of CoRe heads for relevance ranking across all models. For Mistral 7B, CoRe-R substantially outperforms QR-R in performance for any number of retrieval heads. With Llama-3.1 8B (Figure 3(b)), QR-R and CoRe-R are comparable in the re-ranking performance for six or fewer number of retrieval heads, but CoRe-R at $8$ heads surpasses QR-R. Interestingly, with Phi-4 (Figure 3(c)), although QR-R and CoRe-R deliver close average nDCG@10 scores with top $8$ retrieval heads ($51.7$ vs. $52.1$), CoRe-R has a very strong re-ranking performance even with one single attention head ($52.1$), outperforming QR-R at all head levels.
5.3 Re-ranking with Layer Pruning
Our empirical study shows that CoRe-R outperforms existing state-of-the-art list-wise re-rankers. Nevertheless, like other attention-based re-rankers, CoRe-R is memory-intensive, as it requires storing full attention score matrices that scale with the number of tokens. Moreover, computing the attention score matrices introduces latency, which can be particularly costly for list-wise re-rankers.
(a) Mistral 7B
(b) Llama-3.1 8B
(c) Phi-4
Figure 4: Distribution of $S_{CoRe}$ for all heads in each model. We choose the temperature $t=0.001$ for Mistral 7B and $t=0.1$ for Llama-3.1 8B and Phi-4. The top CoRe heads are concentrated mostly in the middle layers of every model.
As illustrated in Figure 4, we observe that the highest scoring CoRe heads are concentrated in the middle layers across all evaluated models. Since attention-based re-ranking does not involve text generation, the final layers, primarily responsible for token generation, are not essential for our task. This observation motivates the use of layer pruning to improve inference efficiency and reduce memory usage. We hypothesize that pruning the majority of the late layers can substantially reduce both memory consumption and computational cost while preserving re-ranking accuracy.
(a) Mistral 7B
(b) Llama-3.1 8B
(c) Phi-4
@@ 5.3 Re-ranking with Layer Pruning
Figure 5: Average latency and re-ranking accuracy on BEIR benchmark. Pruning $50\%$ of the model’s layers does not impact the re-ranking performance while saving $20\%$ inference time.
Figure 6: Peak GPU memory usage on BEIR datasets. Pruning $50\%$ of the model’s layers saves $40\%$ memory usage.
@@
Figure 5 reports the average re-ranking accuracy and latency on BEIR under different levels of layer pruning, while Figure 6 shows the corresponding peak GPU memory usage. The experiment is done on a single GPU H100 96GB. Results show that pruning up to $50\%$ of the late layers preserves re-ranking performance while reducing GPU memory usage by approximately $40\%$ and latency by $20\%$. Beyond this point, performance begins to degrade: pruning $60\%$ of the layers leads to a noticeable decline in the re-ranking performance, underscoring the central role of the CoRe heads in the middle layers. Nevertheless, even with $60\%$ layers pruned, CoRe-R continues to outperform the ICR baseline across all models. This indicates that the remaining CoRe heads in earlier layers [^1] Specifically layers 88-$12$ for Mistral 7B and Llama-3.1 8B, and layers 1212-$16$ for Phi-4 carry sufficiently strong and discriminating signals, surpassing the noisy aggregation of all heads in ICR.
5.4 Impact of Head Detection Strategies
@@
In this subsection, we experiment with different set of retrieval heads, including the original retrieval heads (Wu et al., 2025) identified via copy-paste mechanism on Needle-in-a-Haystack task. We refer to these heads as NIAH heads. For Llama-3.1 8B, none of the top $8$ CoRe heads overlaps with the top $8$ NIAH heads, and only $3$ overlap with the top $8$ QR heads. For Mistral 7B, top $8$ CoRe heads overlap with $2$ QR heads and $2$ NIAH heads. In addition, CoRe heads concentrate in lower layers than both QR heads and NIAH heads. We provide details of the head location in Appendix B.
@@
Table 3 compares CoRe-R with the attention-based re-rankers using the NIAH heads (NIAH-R), QR heads (detected without calibration) and the original QR heads (detected with calibration), for Llama-3.1 8B. Results for ICR, QR-R (without calibration) and CoRe-R are duplicated from Table 1 for reference. As NIAH retrieval heads are limited to copy-paste operations, NIAH-R underperforms ICR on average, showing gains only on a few datasets such as Climate-FEVER and ArguAna, while showing worse performance on many other datasets. On average, the QR-R with calibration improves over ICR by $0.7$ points and QR-R without calibration improves by $1.1$ points, indicating that contextual calibration is not essential for the head detection process. In contrast, CoRe-R shows the largest gain of $1.9$ points, yielding the highest overall nDCG@10 among all baselines.
Overall, attention-based re-rankers achieve strong re-ranking performance across different models, with CoRe-R consistently providing the best nDCG@10 scores. However, there remain cases where all attention-based re-rankers underperform the retrieval baseline (see Tables 1 and 3). Notably, on Climate-FEVER and ArguAna, which often require retrieving counter-evidence, attention-based approaches lab behind the encoder retrieval baseline. Similarly, on the duplicated retrieval dataset Quora, performance also declines. We attribute this to the use of a universal instruction prompt across all datasets, which may not align with dataset-specific task objectives. Future work could explore dataset specific prompt designs to further improve re-ranking performance (see Appendix C.2 for an example on Quora dataset).
@@
Table 3: nDCG@10 on BEIR benchmark for Llama-3.1 8B.
Dataset Retriever ICR NIAH-R QR-R QR-R CoRe-R
Baseline (w/ calib) (w/o calib)
Covid 63.1 76.8 73.7 75.2 75.7 75.7
NFCorpus 33.7 35.0 34.2 35.7 36.4 36.7
DBPedia 36.0 38.7 37.1 39.2 39.0 39.4
SciFact 71.3 74.7 74.2 74.4 75.9 75.1
SciDocs 22.5 18.5 19.2 19.2 19.7 20.2
FiQA 36.9 41.3 40.8 43.4 44.0 44.1
NQ 51.6 60.8 57.5 61.4 63.0 63.2
FEVER 85.5 89.2 88.1 88.5 87.4 88.4
Climate-FEVER 30.3 22.5 23.1 22.8 22.9 22.7
HotpotQA 62.9 73.7 72.0 73.5 73.6 73.8
Touche 24.0 26.1 25.5 26.3 26.3 26.4
MSMARCO 30.7 32.9 31.6 34.7 33.9 34.9
Quora 86.7 78.7 80.1 74.5 74.6 82.3
ArguAna 56.4 42.5 51.7 51.2 54.2 55.0
CQADupstack 44.3 41.7 41.4 43.2 43.2 43.9
Average 49.1 50.2 50.0 50.9 51.3 52.1

### v2

@@ 5.2 Results
Table 1: nDCG@10 on the BEIR benchmark top-$60$.
@@
TREC-COVID 63.1 71.5 74.3 74.6 76.9 77.5 77.9 73.8 77.0 77.0 73.7 71.7 71.9
NFCorpus 33.7 32.7 34.3 34.8 34.5 35.5 36.8 27.2 34.6 34.8 32.4 33.0 33.0
DBPedia 36.0 35.3 36.9 37.5 38.6 39.3 39.8 36.4 40.0 40.3 35.9 35.3 36.6
SciFact 71.3 72.1 71.3 73.6 74.3 74.5 75.3 65.2 70.3 70.9 75.1 74.2 74.5
SciDocs 22.5 16.7 17.1 19.3 18.5 18.6 19.6 17.5 20.7 20.7 19.4 19.5 19.6
FiQA 36.9 37.0 39.6 41.6 42.0 42.7 44.2 36.0 42.8 43.0 41.1 40.9 41.9
NQ 51.6 53.1 54.9 56.8 61.0 63.3 63.2 56.9 63.1 63.7 59.5 57.9 60.5
FEVER 85.5 87.4 87.0 88.4 89.1 88.2 88.1 89.0 89.6 89.6 88.4 88.0 86.6
Climate-FEVER 30.3 21.4 20.8 23.2 22.3 21.7 22.6 22.4 24.9 24.5 21.9 21.1 21.0
HotpotQA 62.9 72.2 72.5 74.2 74.9 74.6 75.0 74.4 75.6 75.7 72.0 73.3 73.6
Touche 24.0 23.5 25.8 27.6 26.9 27.4 28.2 19.2 25.3 25.6 23.5 25.2 25.6
MSMARCO 30.7 28.8 31.1 31.0 32.8 34.5 34.8 30.1 34.5 34.7 30.7 30.4 31.8
Quora 86.7 81.2 71.4 85.1 78.0 72.9 81.5 75.2 70.8 80.5 79.5 70.1 74.5
ArguAna 56.4 46.6 51.2 53.3 43.0 50.5 55.1 40.7 49.8 51.7 54.7 54.6 56.7
CQADupstack 44.3 38.3 40.3 41.2 41.7 43.1 43.4 37.3 42.9 43.5 42.3 43.0 43.6
Average 49.1 47.8 48.6 50.8 50.2 50.9 52.4 46.7 50.8 51.7 50.0 49.2 50.1
Table 2: nDCG@10 on the MLDR datasets top-$40$.
@@
Re-ranking performance on BEIR.
Table 1 reports re-ranking results on the BEIR benchmark. Across all four models, CoRe-R achieves the strongest performance, consistently outperforming both ICR and QR-R. Aggregating attention from all heads in ICR yields the weakest results, often even below the retriever baseline for Mistral 7B and Phi-4, while CoRe-R provides clear and uniform improvements over both the retriever and ICR baselines. Relative to QR-R, CoRe-R shows the largest gains on Mistral 7B, improving average nDCG@10 by +2.2+2.2 over QR-R and +3.0+3.0 over ICR, and on Llama-3.1 8B by +1.5+1.5 and +2.2+2.2 points respectively. Although QR-R remains competitive on the stronger Phi-4 model, its performance varies across datasets, whereas CoRe-R delivers consistently robust gains. Overall, these results demonstrate that CoRe-R reliably isolates informative attention heads across models and datasets, establishing it as a strong state-of-the-art list-wise re-ranker.
We observe that CoRe-R shows remarkable improvement on Quora compared to all of the baselines in Table 1. Quora is a duplicate-question retrieval task with a high proportion of hard negatives, i.e., irrelevant documents that closely resemble the gold answer. In such cases, the contrastive metric shines as it explicitly rewards heads that emphasize the positive document while suppressing near-duplicate. As a result, CoRe-R performs extremely well, while QR-R shows a drastic decline across all models. This aligns well with our theoretical analysis in Proposition 4.2, demonstrating that CoRe heads is more optimal. This furthur suggests that contrastive head selection is particularly valuable for domains with semantically dense or near-duplicate content (e.g., FAQ retrieval, paraphrase search, code retrieval), highlighting an advantage of CoRe-R not captured by nDCG@10 alone.
Cross-lingual and multi-hop generalization.
Table 2 reports the nDCG@10 on the MLDR datasets across three language models. We exclude Phi-4 as the model is not intended to support multilingual use (Microsoft Research, 2024). To ensure compatibility, we evaluate only languages that are supported by all three models. With the same set of CoRe heads detected using a subset of the NQ dataset, CoRe-R shows major improvement in the re-ranking accuracy over all baselines. For Mistral 7B, both ICR and QR-R underperform the retriever baseline on average, while CoRe-R achieves an average improvement of 2.8 points. For Llama-3.1 8B and Granite-3.2 8B, CoRe-R attains the best nDCG@10 scores on all languages with an exception of German for Granite. Figure 5 depicts the Recall@5 score on the multi-hop dataset MuSiQue, and CoRe-R continues to deliver the highest accuracy across all models with an average gain of $2.5$ points over ICR and $1.2$ points over QR-R (see Appendix C.4 for more results). Overall, CoRe-R demonstrates consistent gains on both cross-lingual and multi-hop tasks, underscoring the strong generalization ability of CoRe heads.
Figure 3: Recall@5 on multi-hop dataset MuSiQue top-$40$.
Figure 4: Average nDCG@10 on BEIR with Mistral 7B.
Figure 5: Average nDCG@10 on BEIR with Llama-3.1 8B.
Long context scalability.
Figure 5 and Figure 5 report the average accuracy on BEIR benchmark under varying context lengths. Across all settings, CoRe-R consistently outperforms both ICR and QR-R. For Mistral 7B, all methods experience noticeable degradation as the context length increases. This trend aligns with prior observations that the Mistral 7B model is not robust under long-context settings, e.g., attempts to extend its context window through long-context fine-tuning have been shown to degrade model quality (Scale AI, 2024). For Llama-3.1 8B, both baselines also show mild degradation at longer context lengths. In contrast, CoRe-R continues to improve as context grows, further widening the performance gap. This behavior highlights the strong scalability of CoRe heads and their ability to effectively leverage longer retrieval contexts.
@@
Figure 6: Average latency and re-ranking accuracy on BEIR benchmark top-$40$. Pruning $50\%$ of the model’s layers does not impact the re-ranking performance while saving $20\%$ inference time.
5.3 Efficient Re-Ranking via Layer Pruning
Our empirical study shows that the highest scoring CoRe heads are concentrated in the middle layers across all evaluated models. Since attention-based re-ranking does not involve text generation, the final layers are not essential for our task. We hypothesize that pruning the majority of the late layers greatly reduces memory consumption and computational cost while preserving re-ranking accuracy.
Figure 7: Peak GPU memory usage on BEIR datasets.
@@ 5.3 Efficient Re-Ranking via Layer Pruning
Figure 6 reports the average re-ranking accuracy and latency on BEIR under different levels of layer pruning, and Figure 7 shows the corresponding peak GPU memory usage. The experiment is done on a single GPU H100 96GB. We further report the pruning experiment on the baselines retrieval heads (Wu et al., 2025) and QR heads in Appendix C.3, and the results show that their performance remains stable only under light pruning (30-40%) and begin to degrade once pruning exceeds 40%. Meanwhile, CoRe-R preserves near identical re-ranking performance under $50\%$ pruning, reducing GPU memory usage by approximately $40\%$ and latency by $20\%$. Beyond this point, CoRe-R performance begins to degrade: pruning $60\%$ of the layers leads to a noticeable decline in the re-ranking performance, underscoring the central role of the CoRe heads in the middle layers. Nevertheless, even with $60\%$ layers pruned, CoRe-R continues to outperform the ICR baseline across all models. This indicates that the remaining CoRe heads in earlier layers [^1] Specifically layers 88-$12$ for Mistral 7B and Llama-3.1 8B, and layers 1212-$16$ for Phi-4 carry sufficiently strong and discriminating signals, surpassing the noisy aggregation of all heads in ICR.
5.4 Comparison of CoRe, QR, and NIAH Retrieval Heads
@@
In this subsection, we discuss the overlapping between different set of retrieval heads and compare their performance on final re-ranking tasks. For Llama-3.1 8B, none of the top $8$ CoRe heads overlaps with the top $8$ NIAH heads, and only $3$ overlap with the top $8$ QR heads. For Mistral 7B, top $8$ CoRe heads overlap with $2$ QR heads and $2$ NIAH heads. In addition, CoRe heads concentrate in lower layers than both QR heads and NIAH heads. We provide details of the head location in Appendix B.1.
@@
Table 3 compares CoRe-R with the attention-based re-rankers using the NIAH heads (NIAH-R) for Llama-3.1 8B. As NIAH retrieval heads are limited to copy-paste operations, NIAH-R underperforms ICR on average, showing gains only on a few datasets such as Climate-FEVER and ArguAna, while showing worse performance on many other datasets. Across all datasets, CoRe-R shows the highest overall nDCG@10 compared to all baselines.
5.5 Ablation Study
Effect of detection dataset.
To evaluate the robustness of CoRe head selection, we repeat the detection procedure using a different dataset, MSMARCO. We found that the resulting CoRe heads are nearly identical to those detected from NQ: for both Mistral 7B and Llama-3.1 8B, 7 out of the 8 selected heads overlap, and the one different head comes close within the top-10 heads (see Appendix C.5). We then performed the re-ranking experiments using these MSMARCO-derived heads, and the results are reported in Appendix C.5. With 7 out of 8 heads overlapping, the re-ranking performance closely matches the results obtained using NQ, demonstrating that CoRe head selection is highly stable and does not depend heavily on the detection dataset.
Effect of temperature.
We perform ablation study on the temperature with different values $t=0.1,0.01,0.001$. Since $t$ directly controls the sharpness of the contrastive scoring metric, different temperatures can change the shape of the score distribution and may affect which heads appear in the top-ranked set. We found that the detected CoRe heads in Llama-3.1 8B are consistent across all temperature values while Mistral 7B highly depends on the temperature with 4 common heads detected out of 8 heads (see Appendix B). We observe that many heads in Mistral attend strongly to both the positive document and hard negatives, thus, higher temperatures $t=0.1,0.01$ were not able to capture high-quality heads in Mistral 7B.
Table 3: nDCG@10 on BEIR benchmark top-$40$ for Llama-3.1 8B.
Dataset Retriever ICR NIAH-R QR-R CoRe-R
Baseline
TREC-COVID 63.1 76.8 73.7 75.7 75.7
NFCorpus 33.7 35.0 34.2 36.4 36.7
DBPedia 36.0 38.7 37.1 39.0 39.4
SciFact 71.3 74.7 74.2 75.9 75.1
SciDocs 22.5 18.5 19.2 19.7 20.2
FiQA 36.9 41.3 40.8 44.0 44.1
NQ 51.6 60.8 57.5 63.0 63.2
FEVER 85.5 89.2 88.1 87.4 88.4
Climate-FEVER 30.3 22.5 23.1 22.9 22.7
HotpotQA 62.9 73.7 72.0 73.6 73.8
Touche 24.0 26.1 25.5 26.3 26.4
MSMARCO 30.7 32.9 31.6 33.9 34.9
Quora 86.7 78.7 80.1 74.6 82.3
ArguAna 56.4 42.5 51.7 54.2 55.0
CQADupstack 44.3 41.7 41.4 43.2 43.9
Average 49.1 50.2 50.0 51.3 52.1
Table 4 shows the re-ranking accuracy per dataset under different temperatures. As expected, the results for Llama are similar across all three temperature values, while Mistral demonstrates higher impact from the different detected head set. While a small amount of temperature tuning is necessary for certain models, the entire tuning process incurs minimal overhead as head detection is a relatively fast one-time process per language model. Once selected, the same CoRe heads generalize reliably across cross-domain tasks, multi-hop retrieval, and long-context settings. This demonstrates that CoRe’s performance is not sensitive to temperature in practice, and the detected heads remain robust across applications.
@@ 5.5 Ablation Study
Table 4: nDCG@10 of CoRe-R on BEIR benchmark top-$40$ with different temperature values.
Dataset Retriever Mistral 7B Llama-3.1 8B
Baseline $0.1$ $0.01$ $0.001$ $0.1$ $0.01$ $0.001$
TREC-COVID 63.1 73.6 73.8 73.8 75.7 76.3 76.3
NFCorpus 33.7 34.9 35.6 35.2 36.7 36.6 36.6
DBPedia 36.0 37.1 37.8 37.5 39.4 39.1 39.1
SciFact 71.3 72.2 73.0 74.7 75.1 75.9 75.9
SciDocs 22.5 17.1 18.4 19.5 20.2 20.0 20.0
FiQA 36.9 39.5 41.2 41.4 44.1 44.3 44.3
NQ 51.6 55.6 57.8 58.1 63.2 63.6 63.6
FEVER 85.5 86.9 87.7 88.7 88.4 88.2 88.2
Climate-FEVER 30.3 21.1 21.9 23.3 22.7 22.9 22.9
HotpotQA 62.9 71.6 72.4 73.4 73.8 73.4 73.4
Touche 24.0 26.6 27.1 26.2 26.4 26.9 26.9
MSMARCO 30.7 31.7 32.3 31.7 34.9 34.8 34.8
Quora 86.7 73.4 81.5 85.2 82.3 74.5 74.5
ArguAna 56.4 50.8 52.0 52.7 55.0 54.5 54.5
CQADupstack 44.3 40.3 41.4 41.6 43.9 43.6 43.6
Average 49.1 48.8 50.3 50.9 52.1 51.6 51.6
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning by improving attention-based retrieval and re-ranking methods. The primary impact of this work is technical in nature, contributing to more effective and potentially more efficient model architectures. The societal implications of this work are consistent with those of machine learning research more broadly. Improvements in retrieval and ranking may benefit a wide range of downstream applications that rely on information access and model inference. At the same time, this work does not introduce new application domains or deployment settings beyond those already well studied in the literature. We do not anticipate any novel ethical concerns beyond those commonly associated with machine learning systems.